# OLMoTrace: Tracing Language Model Outputs Back to Trillions of Training Tokens

### v1

@@ 1 Introduction
Tracing the outputs of language models (LMs) back to their training data is an important problem. As LMs gain adoption in higher-stakes scenarios, it is critical to understand why they generate certain responses. However, these modern LMs are trained on massive text corpora with trillions of tokens, which are often proprietary. Fully open LMs (e.g., OLMo; OLMo et al. 2024) enable access to the training data, but existing behavior tracing methods (Koh and Liang, 2017; Khalifa et al., 2024; Huang et al., 2024) have not been scaled to work with this multi-trillion-token setting due to their heavy computation needs.
@@ 3 The Inference Pipeline
During development, we tried keeping the longest spans instead of those with smallest span unigram probability. However, we found that ranking with the span length metric leads to worse relevance level on documents retrieved from the filtered spans (see App. §C and Table 3), and thus we favored the span unigram probability metric. We chose unigram over bigram or trigram because pre-caching them is non-trivial.
@@
To prioritize showing the most relevant documents, in the document panel we rank all documents by a BM25 score in descending order. The per-document BM25 score is computed by treating the collection of retrieved documents as a “corpus”, and the concatenation of user prompt and LM response as the “query”. [^4] We use the implementation in https://github.com/dorianbrown/rank_bm25 We use this BM25 score because it has fairly high agreement with human judgment on topical relevance (§4.2), and can be quickly computed using CPUs. Subsequently, we bucket the relevance scores into three levels – high relevance, medium relevance, and low relevance – and display a colored sidebar on each document to represent its relevance level. High relevance are highlighted with the most saturated color, and low relevance with the least saturated color. We also apply this differential coloring on span highlights: a span’s relevance level is computed as the maximum relevance level among documents enclosing the span. As a result, users are more likely to find highly relevant documents for spans highlighted with the most saturated color.
@@ 3.1 Fast Span Computation
Infini-gram is a text search engine. It supports efficiently counting text queries and retrieving matching documents in massive text corpora with trillions of tokens. To make operations fast, infini-gram indexes text corpora with the suffix array (SA) data structure, and at inference time keeps the huge index files on low-latency SSD disks to avoid loading them into RAM. For OLMoTrace, we build an infini-gram index on the tokenized version of the LMs’ training data (using the Llama-2 tokenizer). On top of this index, in this work we devise a novel parallel algorithm to compute maximal matching spans with low compute latency (Figure 3); we discuss this algorithm and its implementation below.
@@
The problem of finding all maximal matching spans can be broken down into two steps: (1) finding the longest matching prefix of each suffix of the LM output; and (2) suppressing the non-maximal spans. This is because starting from each position, there can be at most one span that is a maximal matching span (if there are two, then one is a subspan of the other and thus is not maximal). Now the first step consists of multiple independent tasks that can be parallelized, and as we will show below, each task can be done with one Find query. Find is a core query operation in infini-gram; it returns the segment of SA that corresponds to all occurring positions of a search term in the text corpus. Since in infini-gram, the processing speed of Find queries is bounded by disk I/O latency and there is a lot of unused throughput, parallelizing these queries can reduce the overall compute latency. In fact, with parallelization, the overall processing speed is bottlenecked by the disk I/O throughput, and thus in our production system we put the index files on high-IOPS SSD disks.
@@
4 Benchmarking and Analysis
@@ 4.1 Benchmarking Inference Latency
4.2 Evaluating Document Relevance
To improve the user experience, OLMoTrace reranks the retrieved documents by relevance to the LM output. We conducted a study to evaluate the relevance level of the top displayed documents. We first composed a rubric for scoring document relevance on a 0–3 Likert scale (App. Table 2, left), and asked a human expert to annotate the top-5 displayed documents for each of the 98 threads used in §4.1 according to this rubric. This human evaluation round was done with OLMoTrace results under a different hyperparameter setting than our final setting, and we later improved the setting under the guidance of LLM-as-a-Judge evaluation. The first document displayed in each thread received an average relevance score of 1.90, and the top-5 documents scored an average of 1.43 (App. Table 3). We then switched to LLM-as-a-Judge evaluation (Zheng et al., 2023) with gpt-4o, and found that it mostly agrees with human evaluation (with a Spearman correlation coefficient of 0.73). LLM-as-a-Judge assigned slightly lower scores overall, with average scores of 1.73 and 1.28 on first and top-5 documents, respectively. We then used LLM-as-a-Judge to guide the tuning of several hyperparameters in OLMoTrace, and our final setting achieved average LLM-as-a-Judge scores of 1.82 on first documents and 1.50 on top-5 documents. See App. §C for additional details on relevance evaluation and hyperparameters tuning.
@@ 5 Case Studies
We envision that researchers and the general public can use OLMoTrace in many ways to understand the behavior of LMs. Below we discuss three example use cases, and we invite the community to explore more use cases.
@@
If the LM states a fact, users may be able to fact-check the statement against its training data. In Figure 4(a), OLMo outputs “The space needle was built for the 1962 World Fair,”. OLMoTrace indicates through the highlight that this span of tokens appears verbatim in the training data and shows the corresponding documents (the screenshot captured one of the 10). For most documents from the pretraining data (like this one), users can click on the “View Document” button and find the URL to the original webpage where this document was crawled. Inspecting the document context and source helps users make a more informed judgment on the factuality of the statement, as words can be misleading out of context, and some web sources may be unreliable.
@@ 6 Related Work
Traditional search engines (e.g., Google) retrieve documents from their web index and may drop results for efficiency or profitability. OLMoTrace faithfully retrieves all matches in an LM’s training data.
@@
OLMoTrace finds lexical, verbatim matches between an LM’s output and its training data. The retrieved documents should not be interpreted as having a causal effect on the LM output, or as supporting evidence or citations for the LM output. In addition, OLMoTrace could make potentially problematic contents in the LM training data more easily exposed, including copyright, PII (personal identifiable information), and toxicity.
@@ Acknowledgements
We would like to thank Alisa Liu, Valentina Pyatkin, Dany Haddad, Shannon Zejiang Shen, Ziqi Ma, and other members of Ai2, H2lab and Xlab for sharing their valuable feedback.

### v2

@@ 1 Introduction
Tracing the outputs of language models (LMs) back to their training data is an important problem. As LMs gain adoption in higher-stakes scenarios, it is critical to understand why they generate certain responses. However, these modern LMs are trained on massive text corpora with trillions of tokens, which are often proprietary. Fully open LMs (e.g., OLMo; OLMo et al. 2024) enable access to the training data, but existing behavior tracing methods (Koh and Liang, 2017; Khalifa et al., 2024; Huang et al., 2024) have not been scaled to work within this multi-trillion-token setting due to their heavy computational needs.
@@ 3 The Inference Pipeline
During development, we tried keeping the longest spans instead of those with smallest span unigram probability. However, we found that ranking with the span length metric leads to worse relevance level on documents retrieved from the filtered spans (see measurement of relevance in App. §C and Table 3), and thus we favored the span unigram probability metric. We chose unigram over bigram or trigram because they computing them (either online or pre-caching) takes a lot of time.
@@
To prioritize showing the most relevant documents, in the document panel we rank all documents by a BM25 score in descending order. The per-document BM25 score is computed by treating the collection of retrieved documents as a “corpus”, and the concatenation of user prompt and LM response as the “query”. [^4] We use the implementation in https://github.com/dorianbrown/rank_bm25 We use this BM25 score because it has fairly high agreement with human judgment on topical relevance (§4), and can be quickly computed using CPUs. Subsequently, we bucket the BM25 scores into three levels – high relevance, medium relevance, and low relevance – and display a colored sidebar on each document to represent its relevance level. High relevance are highlighted with the most saturated color, and low relevance with the least saturated color. We also apply this differential coloring on span highlights: a span’s relevance level is computed as the maximum relevance level among documents enclosing the span. As a result, users are more likely to find highly relevant documents for spans highlighted with the most saturated color.
@@ 3.1 Fast Span Computation
Infini-gram is a text search engine. It supports efficiently counting text queries and retrieving matching documents in massive text corpora with trillions of tokens. To make operations fast, infini-gram indexes text corpora with the suffix array (SA) data structure, and at inference time keeps the huge index files on low-latency SSD disks to avoid loading them into RAM. For OLMoTrace, we build an infini-gram index on the tokenized version of the LMs’ training data (using the Llama-2 tokenizer). On top of this index, in this work we devise a novel parallel algorithm to compute maximal matching spans with low compute latency (Figure 3 and Algorithm 1); we discuss this algorithm and its implementation below.
Algorithm 1 Compute maximal matching spans.
Input Model output $S_{1:L}$ (tokenized), training text corpus $T_{1:N}$ (tokenized) and its suffix array $A_{1:N}$
procedure GetMaximalMatchingSpans($S,T,A$)
     spans [$\leftarrow$[]← [ ]
     for $b$ = 1, …, $L$ do $\triangleright$ execute in parallel
         if $S_{b}$ is a begin-of-word token then
              $len\leftarrow$ GetLongestPrefixLen($S_{b:L},T,A$)
              spans $\leftarrow$ spans + $[(b,b+len)]$
     return SuppressNonMaximalSpans(spans)
procedure GetLongestPrefixLen($s,T,A$)
     $(l,r)\leftarrow$ Find($s,T,A$) $\triangleright$ an infini-gram query
     if $l\neq r$ then $\triangleright$ non-empty segment, $s$ is found in $T$
         $len\leftarrow|s|$
     $\triangleright$ empty segment, $s$ is not found in $T$
         $len1\leftarrow$ LongestPrefixLen($s,T_{A[l]:}$)
         $len2\leftarrow$ LongestPrefixLen($s,T_{A[l+1]:}$)
         $len\leftarrow\max(len1,len2)$
     while $s_{:len-1}$ contains a delimiter token OR $s_{len+1}$ is not a begin-of-word token do
         $len\leftarrow len-1$
     return $len$
procedure SuppressNonMaximalSpans(spans)
     sort spans by beginning position in ascending order
     newspans [$\leftarrow$[]← [ ]
     maxend $\leftarrow$ 0
     for $(b,e)$ in spans do
         if maxend $<e$ then
              maxend $\leftarrow e$
              newspans $\leftarrow$ newspans + $[(b,e)]$
     return newspans
@@
The problem of finding all maximal matching spans can be broken down into two steps: (1) finding the longest matching prefix of each suffix of the LM output; and (2) suppressing the non-maximal spans. This is because starting from each position, there can be at most one span that is a maximal matching span (if there are two, then one is a subspan of the other and thus is not maximal). The first step consists of multiple independent tasks that can be parallelized, and as we will show below, each task can be done with one Find query. Find is a core query operation in infini-gram; it returns the segment of SA that corresponds to all occurring positions of a search term in the text corpus. Since in infini-gram, the processing speed of Find queries is bounded by disk I/O latency and there is a lot of unused throughput, parallelizing these queries can reduce the overall compute latency. In fact, with parallelization, the overall processing speed is bottlenecked by the disk I/O throughput, and thus in our production system we store the index files on high-IOPS SSD disks.
@@
Figure 4: Statistics of spans and documents outputted by OLMoTrace. Left: The spans selected are relatively long, with a mean length of 10.4 tokens. Middle: The max attainable BM25 relevance score of a document is roughly proportional to the response length, so we make the score thresholds proportional to the response length. Right: The relevance score of a span is the max score among its corresponding documents.
@@ 3.2 Benchmarking Inference Latency
4 Analyses
We analyze the some properties of the spans and documents outputted by OLMoTrace, using the same 98 conversations as in §3.2.
Length of spans.
We report the length of spans given after step 2 (filtering for long and unique spans, before merging). The spans have a mean length of 10.4 tokens and a median of 10 tokens. Figure 4 (left) shows the distribution of span lengths. This tells us that there are many long pieces of text shared between the LM output and its training data, which are revealed by OLMoTrace.
Relevance score of documents.
To improve user experience, OLMoTrace reranks the retrieved documents by relevance to the LM output using BM25. We found that the maximum attainable BM25 score is roughly capped by 0.18 times the number of characters in the LM output (Figure 4, middle), so we normalize the BM25 scores by this coefficient. After normalization, we bucket the scores as follows: $\geq 0.7$ is high relevance, between 0.5 and 0.7 is medium relevance, and $<0.5$ is low relevance. We empirically found these thresholds to be aligned with human expectations, and this puts 14% of documents as high relevance. We use the same normalization and thresholds for span scores (Figure 4, right), rendering 19% of spans as high relevance.
Validating the relevance rankings.
We conducted a study to evaluate the relevance level of the top displayed documents to the LM output. We first composed a rubric for scoring document relevance on a 0–3 scale (App. Table 2, left), and asked a human expert to annotate the top-5 displayed documents for each conversation according to this rubric. This human evaluation round was done with OLMoTrace results under a different hyperparameter setting than our final setting, and we later improved the setting under the guidance of LLM-as-a-Judge evaluation. The first document displayed in each conversation received an average relevance score of 1.90 (roughly meaning “being on the same topic as the LM output), and the top-5 documents scored an average of 1.43 (App. Table 3). We then switched to LLM-as-a-Judge evaluation (Zheng et al., 2023) with gpt-4o, and found that it mostly agrees with human evaluation (with a Spearman correlation coefficient of 0.73). LLM-as-a-Judge assigned slightly lower scores overall, with average scores of 1.73 and 1.28 on first and top-5 documents, respectively. We then used LLM-as-a-Judge to guide the tuning of several hyperparameters in OLMoTrace, and our final setting achieved average LLM-as-a-Judge scores of 1.82 on first documents and 1.50 on top-5 documents. See App. §C for additional details on relevance evaluation and hyperparameters tuning.
Training stage of retrieved documents.
Among the retrieved documents, we found the vast majority (96.7%) belong to the pre-training data, 0.9% belong to the mid-training data, and 2.4% to the post-training data. Among post-training, 0.9% are from the SFT data, 1.5% are from the DPO data, and none are from the RLVR data. We note that this distribution heavily depends on the topic of the conversation: for example, a math-heavy LM output may result in more documents retrieved from SFT and RLVR datasets.
@@ 5 Case Studies
We envision that researchers and the general public can use OLMoTrace in many ways to understand the behavior of LMs. Below we discuss three example use cases, and we invite the community to explore additional ones.
@@
If the LM states a fact, users may be able to fact-check the statement against its training data. In Figure 5(a), OLMo outputs “The space needle was built for the 1962 World Fair,”. OLMoTrace highlights this span of tokens as it appears verbatim in the training data and shows the corresponding documents (the screenshot captured one of the ten documents). For most documents from the pretraining data (like this one), users can click on the “View Document” button and find the URL to the original webpage where this document was crawled.
We note that inspecting the document context and source can help users make a more informed judgment on the factuality of the statement, as words can be misleading out of context, and some web sources may be unreliable.
@@ 6 Related Work
Traditional search engines (e.g., Google) retrieve documents from their web index. OLMoTrace retrieves matches in an LM’s training data, which is more suitable to use for understanding the data origin of LM behaviors.
@@
OLMoTrace finds lexical, verbatim matches between an LM’s output and its training data. The retrieved documents should not be interpreted as having a causal effect on the LM output, or as supporting evidence or citations for the LM output.
Mitigating social and legal risks.
OLMoTrace can make potentially problematic contents in the LM training data more easily exposed. We conducted an internal red-teaming effort and implemented mitigation measures based on the findings. We focused on three aspects: copyright, PII (personal identifiable information), and toxicity. For copyright, we were able to make OLMoTrace show documents with news articles or song lyrics, while we did not see any copyrighted book; we offer a takedown request form for copyright holders to fill out in case they identify documents infringing their copyright, and we implemented an efficient way to take down documents in the infini-gram engine so that we don’t need to re-index the full training data. For PII, we were unable to find any PII data in OLMoTrace results, and we implemented a regex-based filter to block documents with PII. For toxicity, text moderation is already implemented in Ai2 Playground to filter user prompts, and we do not add further filtering.
@@ Acknowledgements
We would like to thank Alisa Liu, Valentina Pyatkin, Dany Haddad, Shannon Zejiang Shen, Ziqi Ma, and other members of Ai2, H2lab and Xlab for sharing their valuable feedback. This work was funded in part by NSF IIS-2044660.